If x 2/3 = 1/4 and x is a whole number ​then find the value of 2x + 3

  1. If x2/3=1/4 and x is a whole number
  • ​then find the value of 2x + 3

Grade:11

2 Answers

Nitesh
6 years ago
From the question x to the power 2/3=1/4 so we get x= 1/4 to the power 3/2. We get x=1/8.Hence 2x+3=2×1/8 + 3=13/4 .So the answer is 13/4.
yathartha gupta
6 years ago
From the question x to the power 2/3=1/4 so we get x= 1/4 to the power 3/2. We get x=1/8.Hence 2x+3=2×1/8 + 3=13/4 .So the answer is 13/4